# Experimental Measurements of the Granular Density of Modes via Impact

### v1

@@ II.2 Piezoelectric Sensors
If these sensors are simply placed into the granular material without any additional circuitry, a voltage could be measured in some proportion to acoustic vibrations. However, this signal would be very weak, possibly not even exceeding the noise threshold. It is therefore necessary to connect the piezoelectric sensors to pre-processing circuitry to amplify the signal and reduce the noise level. The circuit that we developed (shown in Figure 3) to accomplish this task consists of two charge amplifiers [10], one connected to each side of the crystal. These charge amplifiers convert the charge on each side of the crystal into a voltage. These two voltages are then fed through a differential amplifier for both amplification and common mode noise rejection. This circuit yields high sensitivity and low noise, ideal for measuring small vibrations. Additionally, the output voltage of the circuit, $V_{out}$ can be converted into the charge, $q$, generated by the piezoelectric with the following relation $q=\frac{V_{out}}{2GC_{f}}$, where $G$ is the gain of the differential amplifier and $C_{f}$ is the feedback capacitor.
@@ II.3 Data Collection Protocol
After this, pressure is applied via the screw to the pressure plate, and the ball is dropped again and the response of the piezoelectrics is again recorded. This procedure is repeated for a total of four pressure states (listed above). Once all the pressure states have been cycled through. The grains are removed, mixed, added back into the experimental apparatus, and the whole procedure is repeated. In total, we perform this procedure on five unique packings for each of our three grain types (15 unique packings in total).
Figure 5: The density of modes for (a) the bidisperse system, (b) the 8 mm monodisperse system, and (c) the 6 mm monodisperse system. The different colors correspond to the different pressures that the system is subjected to.
Figure 5: The density of modes for (a) the bidisperse system, (b) the 8 mm monodisperse system, and (c) the 6 mm monodisperse system. The different colors correspond to the different pressures that the system is subjected to.
Figure 6: (a) The DOM for the bidisperse packing zoomed in on the frequencies between 10 Hz and 1 kHz. The location of the primary peaks are labeled with a plus sign and the secondary peak with a star symbol. (b) The frequency of the primary peak with the inset showing the corresponding wavelengths for the 6 mm beads. (c) The frequency of the secondary peak with the inset showing the corresponding wavelengths for the 6 mm beads.
@@ II.4 Sound Speeds
In order to be able to relate specific frequencies to relevant length scales, we need to know the speed of sound in our granular material. This was accomplished for the 6 mm packing by measuring the time difference between the first arrival of sound (time of flight) at two piezoelectric sensors separated from each other by 10 cm. For these sound speed measurements, one piezoelectric sensor is placed near the top of the system and the other is placed 10 cm below this. A pulse of sound is injected at the top of the system and the time at which the signal first arrives at each sensor is found. From this time of flight information, a sound speed for the material is found. The grains are then agitated to put the material in a new state and the sound speed measurement is repeated for a total of 20 unique grain configurations. A histogram of these speeds is shown in Figure 4. We find an average speed of $193\pm 16$ m/s (standard error).
@@ III Results and Analysis
First, we recognize that the charge generated on the piezoelectric is proportional to the strain on the crystal, that is $q\propto x$, where $x$ is the change in thickness of the crystal. From this, we can see from both an electrical and mechanical perspective that the energy imparted to the crystal from the granular material is proportional to $q^{2}$. The electrical perspective treats the piezoelectric crystal as a capacitor, and the energy of any capacitor proportional to $q^{2}$. The mechanical perspective treats the crystal as a spring with the energy of a spring proportional to $x^{2}$ which for a piezoelectric is also proportional to $q^{2}$.
Second, since the piezoelectric sensors are only in contact with the granular material, all of the energy imparted to the piezoelectric must come from the kinetic energy of its neighboring grains. Since the kinetic energy of the grains is proportional to $v^{2}$ and the energy of the piezoectric crystal is proportional to $q^{2}$, it follows that $v\propto q$ where the constant of proportionality is some combination of material properties of the grains and the crystal, for example Young’s modulus, geometry, piezoelectric charge coefficient, etc. The fact that $q$ does not have the precise conversion factor to exactly find $v$ in proper units is not necessary for computing the DOM since Equation 1 simply needs to know the relative amplitudes of $v$ at each frequency, and then the normalization of the DOM will happen automatically as a consequence of Equation 1.
@@
Now that the velocity of the particles around the sensors can be deduced, the rest of the calculation follows straightforwardly. For each particle type, i.e. 6 mm monodisperse, 8 mm monodisperse, or bidisperse, and at each pressure the velocity autocorrelation function, $C_{v}(t)$, is found by finding the average charge autocorrelation function for all the sensors. Then, as described in Equation 2, the FFT of $C_{v}$ gives the density of modes. Now that we have obtained the density of modes, a final smoothing step is taken by applying a median filter with a 15 Hz window to the density of modes. Compared to the 25 kHz data range, a 15 Hz filter window does not substantially change the density of modes other than a slight flattening of some of the peaks, but it does smooth some of the noise fluctuations seen in the data. The results of these calculations are shown Figure 6. These results are reported in the frequency range of 10 Hz to 25 kHz. 10 Hz is the lowest frequency that we report due to the fact that the impactor ball has a bouncing period of approximately 10 Hz and the low frequency 3dB point of our piezoelectric circuitry is 16 Hz. The high frequency cutoff of 25 kHz is set by the Nyquist frequency.
Some features that are immediately apparent from the density of modes in Figure 6 are the excess low frequency modes compared to the Debye scaling as well as a drop off in the number of modes around 10 kHz. We also see several peaks in the approximately 10–1000 Hz range; for a better visual example, Figure 5(a) is a zoomed in graph of the density of modes in this frequency range. Looking at Figure 5(a), there are two peaks of particular interest that shift with system pressure. The first peak is the overall maximum peak in the system, call this the primary peak, the second peak is around 100 Hz, call this the secondary peak, which also shifts with system pressure. The change in the frequency, $\Delta f$, at which each of these peaks occurs relative to the low pressure peak is found and plotted as a function of system pressure in Figures 5(b) and 5(c), and it is seen that the peaks move to higher frequency with increasing system pressure.
In order to explore these two peaks in the density of modes, we calculate the wavelength corresponding to these frequencies using the sound speed and the 6 mm beads. However, as was shown in [12] the sound speed in a granular material is not constant with respect to pressure, and for spheres following the Hertzian contact law [9], the sound speed should increase with pressure as $P^{1/6}$ [12]. Applying the $P^{1/6}$ correction to the sound speed, we calculate the wavelengths corresponding to the primary and secondary peaks for the 6 mm beads and show the results in the inset in Figures 5(b), and 5(c).
@@ IV Discussion
One of the first things seen in Figure 6 is that around 10 kHz the density of modes falls off and then levels off to a minimum value. In the Debye model of the density of states, the highest phonon mode has a wavelength of twice the inter-atomic spacing. If we attempt to translate the concept of inter-atomic spacing to a granular system, we might expect the shortest wavelength to be twice the grain diameter. Using the sound speed for our 6 mm monodisperse system leads to a frequency of $16\pm 1.3$ kHz which is within a factor of two from the experimental value. We also see that at the lowest pressure, the density of modes reaches its high frequency cutoff at a lower frequency than when the system is under higher pressure.
The next observation is that the primary peak, shown in Figures 5(a) and 5(b), increases with increasing pressure, and the corresponding wavelength for the 6 mm beads ranges from approximately 50 cm to 60 cm (see inset in Figure 5(b)). These wavelengths are approximately twice the system width and may correspond to a standing wave of the system size, which would explain its dominance in the density of modes. An interesting method to validate this hypothesis would be to create a system that had different dimensions for the length, width, and height of the box and see if multiple peaks in the density of modes emerge.
The secondary peak in Figure 5(c) is also seen to increase with system pressure, and the corresponding wavelengths of the 6 mm beads are shown in the inset in Figure 5(c). The wavelengths of these peaks corresponds to approximately 2 m which is much larger than the system size. The wavelengths also seem to increase with pressure in contrast to the behavior of the primary peak. The fact that these peaks do change with system pressure gives us confidence that they are real features of the granular material and not spurious electrical noise or artifacts. Additionally, there are several other peaks that also appear to scale with system pressure around the 200 Hz mark. These peaks are also soft modes of the system and deserve future investigation.
Another peak of note in the density of modes is the peak seen at 60 Hz. This peak is slightly flattened due to the median filter and the logarithmic scaling of the figure but is well centered at 60 Hz regardless of system pressure. This peak is almost certainly an electrical artifact of the 60 Hz mains in the building’s electrical wiring.
@@
An additional point of investigation in this study is the effect of different grain size as well as monodisperse packings versus bidisperse packings. Looking at Figure 7, we see that while there are some differences in the density of modes between the three systems, it is not a dramatic difference, with general agreement across the main features of the density of modes. This is not particularly surprising given that the grains are all made of the same material and of the same shape (spherical). Additionally, while [13] did see a difference between monodisperse packings and bidisperse packing in 2D due to the crystallization of the 2D packing, no such regular crystallization occurs in 3D.
While there are broad similarities in the density of modes across the different packings, they are not identical. This can be seen in Figures 5(b) and 5(c). In particular, in both plots there is a difference between the 6 mm and 8 mm packings. Curiously, in Figure 5(b) the bidisperse packing closely resembles the 8 mm packing, and then in Figure 5(c) the bidisperse packing switches and closely resembles the 6 mm packing.
Finally, one of the strengths of this work is the novel approach to exciting the vibrational modes with a simple impact event. In addition to being experimentally simple, an impact has a particular intrinsic advantage over other excitation methods. Namely, while the impactor is striking the granular material, the system can be thought of as a damped, driven oscillator, and since the impact event is very short, transients will be important to the response. For damped, driven oscillators, the transient terms oscillate at the natural frequency of the system and not the driving frequency, which is exactly what is needed for our experiments. Then, after the impact event, the driving force is no longer present and the system behaves more like a damped oscillator, essentially ringing at the natural frequencies of the material.
To fully take advantage of this technique it is helpful to make the impact event as short as possible. If the impactor colliding with the impactor plate is thought of as an object colliding with a spring, then the period of that spring’s oscillation would correspond to the duration of the collision. Using this model for the impact time, a short impact can be accomplished with a light, but hard and stiff impactor such as the small, hardened steal ball bearing used in this study. If higher energy is needed from the impactor, it is best accomplished by increasing the speed of the impactor rather than its mass so that the impact time remains small. Finally, so long as the energy of the impactor does not cause the grains to permanently change their position, the energy of the impactor should not influence the $D(\omega)$ of an individual packing, provided there is enough energy to excite the same number of modes with different energies. However, if data from multiple packing are averaged together, as was done in this study, the energy of the impactor should be kept very consistent as packings with higher energy impacts will be weighted more heavily in Equation 1.
@@ V Conclusion
In this work, we have successfully measured the granular density of modes in three dimensions for a packing of spherical grains. Our measurements of the density of modes displays an excess of low frequency modes as the jamming transition is approached. We also identify several interesting peaks that shift with pressure and have length scales at, or longer than, the system size. Future work might vary the dimensions of the granular system in order to see if these peaks shift or split into multiple peaks as the symmetry of the system size is broken. Additionally, we see the density of modes fall off at high frequency in a manner reminiscent of the Debye frequency; that is, the density of states falls off at a wavelength close to twice the particle size. All of these results continue to demonstrate the utility in borrowing solid state physics techniques and applying them to granular systems. Finally, our novel approach of exciting the grains with a simple impact, increases experimental accessibility and will allow for measurements of the granular density of modes to become more commonplace.

### v2

@@ II.3 Data collection protocol
After this, pressure is applied via the screw to the pressure plate, and the ball is dropped again and the response of the piezoelectrics is again recorded. This procedure is repeated for a total of four pressure states (listed above). Once all the pressure states have been cycled through, the grains are removed, mixed, added back into the experimental apparatus, and the whole procedure is repeated. We measure the density of modes as the pressure on the system is cycled from low to high, since we find these results are not qualitatively different from cycling the pressure on the system from high to low. In total, we perform this procedure on five unique packings for each of our three grain types (15 unique packings in total).
Figure 6: The density of modes for (a) the bidisperse system, (b) the 8 mm monodisperse system, and (c) the 6 mm monodisperse system. The different colors correspond to the different pressures that the system is subjected to. The dashed line is the power spectrum of the noise fit from Figure 4. The vertical units of the noise power spectrum are arbitrary in order for the noise spectrum to roughly, vertically align with the high frequency noise in the density of modes. The solid black line is $\propto f^{2}$ for comparison to Debye scaling.
@@ II.4 Sound speeds
In order to be able to relate specific frequencies to relevant length scales, we need to know the speed of sound in our granular material. This was accomplished for the 6 mm packing by measuring the time difference between the first arrival of sound (time of flight) at two piezoelectric sensors separated from each other by 10 cm. For these sound speed measurements, one piezoelectric sensor is placed near the top of the system and the other is placed 10 cm below this. A pulse of sound is injected at the top of the system and the time at which the signal first arrives at each sensor is found. From this time of flight information, a sound speed for the material is found. The grains are then agitated to put the material in a new state and the sound speed measurement is repeated for a total of 20 unique grain configurations. A histogram of these speeds is shown in Figure 5.
Figure 7: (a) The DOM for the bidisperse packing zoomed in on the frequencies between 10 Hz and 1 kHz. The location of the primary peaks are labeled with a plus sign and the secondary peak with a star symbol. (b) The $\Delta f$ of the primary peak as a function of pressure. (c) The $\Delta f$ of the secondary peak as a function of pressure.
We find an average speed of $193\pm 16$ m/s (standard error). This speed of sound is significantly slower than the bulk sound speed, $c=\sqrt{\frac{E}{\rho}}$, of 1,200 m/s. This discrepancy with the bulk speed is consistent with previous experiments [39, 40, 41] that found the granular speed of sound for a granular material composed of unconfined glass beads to be 280 m/s compared to a bulk speed of 4,000 m/s. Additional experiments on a 2D granular material [42] found a wide range of sound speeds within a given packing due to the fact that the speed of sound varies with the strength of the force chain along which the sound propagates.
Additionally, simulations of the sound speed from the time of flight [43] find a pressure dependence of $c\propto P^{1/6}$. In their simulations, they report pressure in units of $E^{*}=\frac{E}{1-\nu^{2}}$ and speeds in units of $c^{*}=\sqrt{\frac{E^{*}}{\rho}}$, which for our 6 mm packing are $E^{*}\approx 3$ GPa and $c^{*}\approx 1700$ m/s. At the pressure at which our sound speeds were measured ($0.7\mu E^{*}$), the simulation finds a sound speed of $2000.12c^{*}=$ 200 m/s in good agreement with our average result of 193 m/s.
@@ III Results and Analysis
First, we recognize that the charge generated on the piezoelectric is proportional to the strain on the crystal, that is $q\propto x$, where $x$ is the change in thickness of the crystal. From this, we can see from both an electrical and mechanical perspective that the energy imparted to the crystal from the granular material is proportional to $q^{2}$. The electrical perspective treats the piezoelectric crystal as a capacitor, and the energy of any capacitor is proportional to $q^{2}$. The mechanical perspective treats the crystal as a spring with the energy of a spring proportional to $x^{2}$ which for a piezoelectric is also proportional to $q^{2}$.
Second, since the piezoelectric sensors are only in contact with the granular material, all of the energy imparted to the piezoelectric must come from the kinetic energy of its neighboring grains. Since the kinetic energy of the grains is proportional to $v^{2}$ and the energy of the piezoelectric crystal is proportional to $q^{2}$, it follows that $v\propto q$ where the constant of proportionality is some combination of material properties of the grains and the crystal, for example Young’s modulus, geometry, piezoelectric charge coefficient, etc. The fact that $q$ does not have the precise conversion factor to exactly find $v$ in proper units is not necessary for computing the DOM since Equation 1 simply needs to know the relative amplitudes of $v$ at each frequency, and then the normalization of the DOM will happen automatically as a consequence of Equation 1.
@@
Now that the velocity of the particles around the sensors can be deduced, the rest of the calculation follows straightforwardly. For each particle type, i.e., 6 mm monodisperse, 8 mm monodisperse, or bidisperse, and at each pressure the velocity autocorrelation function, $C_{v}(t)$, is found by finding the average charge autocorrelation function for all the sensors. Then, as described in Equation 4, the FFT of $C_{v}$ gives the density of modes. Now that we have obtained the density of modes, a final smoothing step is taken by applying a median filter with a 15 Hz window to the density of modes. Compared to the 25 kHz data range, a 15 Hz filter window does not substantially change the density of modes other than a slight flattening of some of the peaks, but it does smooth some of the noise fluctuations seen in the data. The results of these calculations are shown Figure 6. These results are reported in the frequency range of 10 Hz to 25 kHz. 10 Hz is the lowest frequency that we report due to the fact that the impactor ball has a bouncing period of approximately 10 Hz and the low frequency 3dB point of our piezoelectric circuitry is 16 Hz. The high frequency cutoff of 25 kHz is set by the Nyquist frequency. As a check that our results are above the noise threshold of the electronics, we also add the power spectrum of the noise found in Figure 4 to the results in Figure 6. We add the power spectrum due to Wiener-Khinchin’s theorem [44] that equates the power spectrum with the Fourier transform of the autocorrelation function.
Some features that are immediately apparent from the density of modes in Figure 6 are the deviation of the low frequency modes from Debye scaling as well as a drop off in the number of modes around 10 kHz. We also see several peaks in the approximately 10–1000 Hz range; for a better visual example, Figure 7(a) is a zoomed in graph of the density of modes in this frequency range. Looking at Figure 7(a), there are two peaks of particular interest that shift with system pressure. The first peak is the overall maximum peak in the system, call this the primary peak, the second peak is around 100 Hz, call this the secondary peak, which also shifts with system pressure. The change in the frequency, $\Delta f$, at which each of these peaks occurs relative to the low pressure peak is found and plotted as a function of system pressure in Figs. 7(b) and 7(c), and it is seen that the peaks move to higher frequency with increasing system pressure.
@@ IV Discussion
From Figure 8, we see that while there are some differences in the density of modes between the three systems, it is not a dramatic difference, with general agreement across the main features of the density of modes. This is not particularly surprising given that the grains are all made of the same material and of the same shape (spherical). Previous experimental measurements in 2D [31] did see a difference between monodisperse packings and bidisperse packing due to the crystallization of the 2D packing. However, in our 3D packings, a bidisperse mixture is not required to break up crystallization as it is in a 2D packing [45, 46]. Given these observations, we will focus the remainder of our discussion on the 6 mm monodisperse packing; this discussion should also broadly apply to the 8 mm and bi-disperse packings.
In order to interpret our results, it is helpful to compare with prior simulations of the density of modes [20, 3, 25, 26]. We begin our comparison to simulations by converting the simulation units used in Ref. [3] into our experimental units; in these simulations, one unit of frequency defined is $\frac{1}{2\pi}\sqrt{\frac{V_{0}}{ma^{2}}}$ where $a$ is the particle diameter, $m$ is the particle mass, and $V_{0}$ comes from the particle interaction potential, which for spheres gives a force law $F=\frac{V_{0}}{a^{5/2}}\delta^{3/2}$, where $\delta$ is the particle overlap. This simulation force law can be equated to the Hertzian contact force law [47] $F=\frac{\sqrt{a}}{3(1-\nu^{2})}E\delta^{3/2}$. We can now write the simulation unit of frequency in terms of our particle properties as $\sqrt{\frac{aE}{12\pi^{2}m(1-\nu^{2})}}$, which for the 6 mm grains is equal to 28 kHz.
Figure 6 shows the density of modes disappearing into the noise around 10 kHz or 0.36 in simulation units, which is a much lower frequency than the more typical simulation roll off frequency of about $2.5$ seen in Ref. [3]. One might at first suspect that our experiment is simply not exciting modes above 10 kHz; however, Figure 4 shows that we are in fact exciting vibrational modes above 10 kHz. Instead, the system does not “ring” as well at these modes and/or these high frequency modes are more strongly dissipated than their low frequency counterparts. Frequency dependent dissipation would be consistent with the viscoelastic nature of ABS plastic [48, 49] and could represent a deviation from the simulated systems that typically model the grains as elastic. Additionally, the viscoelastic damping of the high frequencies may also account for the more gradual falling off of the high frequency side of the density of modes compared to a very steep drop off seen in simulations. While our short impact time combined with a viscoelastic system may limit the highest frequencies accessible for study, much of the information related to jamming occurs in the lower frequency parts of the density of states where we have good resolution down to approximately 0.0036 in simulation units.
The primary peak, shown in Figs. 7(a) and 7(b) increases with increasing pressure. To investigate this further, these frequencies should be correlated to a length scale via the speed of sound. However, this is not a straightforward endeavour as the speed of sound does not have a consistent value as shown in Figure 5. If the average value of the speed of sound is used, then the primary peak at the lowest pressure has a wavelength of 60 cm which is very close to twice the system size and may correspond to a standing wave across the system. An interesting method to validate this hypothesis would be to create a system that had different dimensions for the length, width, and height of the box and see if multiple peaks in the density of modes emerge. Additionally, the viscoelastic nature of the ABS plastic may also serve to accentuate the primary peak, since without the frequency dependent damping, the density of states may look more flattened at higher frequencies instead of the gradual roll off discussed above. Nonetheless, this primary peak must represent a dominant mode within the system even if its prominence is raised due to viscoelasticity as compared to simulations.
The secondary peak in Figure 7(c) is also seen to increase with system pressure. Using the average sound speed, the wavelength of the secondary peak in the lowest pressure system is 200 cm, which is longer than the system size. However, these modes may not actually correspond to wavelengths longer than twice the system size as the speed of sound varies considerably. The slowest speed of sound that we measured in Figure 5 is 56 m/s, and if this sound speed is used, the wavelength would be 58 cm which is again close to twice the system size. Within a given packing the sound speed is going to vary depending on the strength of the force chain on which the sound propagates [42]. Given that force distribution in a granular material is dominated by a majority of low force particles [50], it is possible that this secondary peak corresponds to these low speed modes.
The observation of these peaks in the density of modes is a departure from what is seen in simulations. However, the fact that these peaks do change with system pressure gives us confidence that they are real features of the granular material and not spurious electrical noise or artifacts. This can be compared to another peak of note in the density of modes seen at 60 Hz. This peak is slightly flattened due to the median filter and the logarithmic scaling of the figure but is well centered at 60 Hz regardless of system pressure. This peak is almost certainly an electrical artifact of the 60 Hz mains in the building’s electrical wiring.
@@ V Conclusion
This work makes two meaningful contributions to the field of granular physics. The first and main contribution is the suggestion and experimental exploration of a technique for exciting the vibrational modes of a granular material via impact in order to measure the granular density of modes. The second contribution of this work is to measure the density of modes in an as yet unmeasured, experimental, three dimensional granular system, ABS plastic spheres; thereby providing an additional experimental data point on the nature of the granular density of modes in real materials.
Our measurements of the density of modes displays an excess of low frequency modes at low pressure consistent with previous work. We also identify several interesting peaks that shift with pressure. Future work might vary the dimensions of the granular system in order to see if these peaks shift or split into multiple peaks as the symmetry of the system size is broken. Additionally, we see the density of modes rolls off on the high frequency side at a slower rate and lower frequency than what is seen in simulation. This deviation from simulation is likely due to the viscoelastic nature of the particles and represents an additional complexity of real materials that is not currently captured in the simulations. In conclusion, our approach of exciting the grains with a simple impact increases experimental accessibility and will allow for measurements of the granular density of modes to become more commonplace.